Physics

A bicycle coasting at 6.0 comes to a 5.0-long, 1.0-high ramp. What is the bicycle's speed as it leaves the top of the ramp?
Answer #1

SOLUTION :


K. E. at the beginning of ramp = 1/2 m v^2 = 1/2 m (6)^2 = 18 m   (N-m) 


P. E at the top of the ramp = m g h = m * 9.8 * 1.0 = 9.8 m    (N-m)


Let the speed at the top of ramp be V m/sec.

So K.E at the top of the ramp = 1/2 m V^2 


So, aș per energy conservation :


Total energy at the beginning of the ramp = Total energy at the top of the ramp

=> 18 m = 9.8 m + 1/2 m V^2 

=> 18 = 9.8 + 1/2 V^2

=> V^2 = (18 - 9.8) * 2 = 16.4 

=> V = sqrt(16.4) = 4.05 m/sec 


So, bicycle speed as it leaves the top of the ramp = 4.05 m/sec (ANSWER).


answered by: Tulsiram Garg